# Modeling the Dynamics of Sub-Millisecond Electroadhesive Engagement and Release Times

### v1

@@ Abstract
Electroadhesion is an electrically controllable switchable adhesive commonly used in soft robots and haptic user interfaces. It can form strong bonds to a wide variety of surfaces at low power consumption. However, electroadhesive clutches in the literature engage to and release from substrates several orders of magnitude slower than a traditional electrostatic model would predict, limiting their usefulness in high-bandwidth applications. We develop a novel electromechanical model for electroadhesion, factoring in polarization dynamics and contact mechanics between the dielectric and substrate. We show in simulation and experimentally how different design parameters affect the engagement and release times of electroadhesive clutches to metallic substrates. In particular, we find that higher drive frequencies and narrower substrate aspect ratios enable significantly faster dynamics. We demonstrate designs with engagement times under 15 µs and release times as low as 875 µs, which are 10$\times$ and 17.1$\times$ faster, respectively, than the best times found in prior literature.
Index Terms:
Electroadhesion, switchable adhesion, dynamics, modeling, actuators
I Introduction
@@
Switchable adhesives can dynamically adhere to and release from contacting surfaces. They are widely used in soft robots, haptic interfaces, and biomedical devices, with applications in robotic grippers [1, 2], wall climbing and perching robots [3, 4], exoskeletons [5, 6], haptic user interfaces [7, 8, 9, 10], and kinesthetic garments [11, 12]. Different switchable adhesion mechanisms can be compared by their shear force capacity when gripping different substrates and by their engagement and release speeds [13]. Mechanisms with high shear force capacities, like granular jamming actuators and shape memory polymers, often require long switching times due to viscoelastic effects, while faster adhesives like liquid crystal elastomers and gecko-inspired adhesives often have caveats like low force capacity or poor adhesion to rough surfaces [14, 15].
@@
Electroadhesion (EA), or the electrostatic attraction between two contacting materials separated by a dielectric and held at a potential difference, has emerged as an effective electrically triggered switchable adhesive. Its scalable manufacturing process, high shear force capacity against a wide range of substrates, relatively fast actuation speeds, low power consumption, and low profile have made EA useful for space- and power-constrained applications. In this paper, we focus on a common implementation, shown in Fig. 1, where one of the contacting materials is a dielectric patterned with interdigitated comb electrodes while the other contacting material can be any conductive or insulating substrate. This form factor, originally developed in 1968 by Krape [16], sacrifices some adhesive force capacity compared to a parallel plate configuration in which both contacting materials have electrodes. However, EA with Interdigitated Co-planar Electrodes (shortened to EA-ICE in this paper) allow all the electronics to be concentrated into a single surface, reducing bulk and enabling the clutch to adhere to small, delicate objects and generic surfaces in the field.
@@
Figure 1: (a) Schematic diagram and (b) picture of an electroadhesive clutch. When a voltage $V$ is applied, the polarization of bound charges in the dielectric amplifies the field strength, attracting the film to the metal substrate. The resulting friction can counteract applied shear loads on the substrate.
@@
Despite its widespread utility, there is a lack of accurate dynamics models for how quickly EA pads can adhere to and release from surfaces. Prior work has shown that the dielectrics themselves can be charged and discharged on the order of microseconds [17, 18]. However, as shown in Fig. 2, every paper we found reported engagement times from hundreds of microseconds to milliseconds and release times ranging from tens of milliseconds to minutes. While some models predicting minute-scale release times have been proposed (although, to our knowledge, never compared numerically against experimental data) [19, 20, 21, 22], experimental results in the microsecond to millisecond range have largely gone unexplained, limiting EA’s scalability to high-bandwidth applications [23, 10]. Slow dynamics also leads to “stick-slip” behavior, where after an applied load exceeds an EA clutch’s force capacity the clutch slips, losing up to 70% of its frictional capacity and requiring tens of milliseconds to re-engage [6, 24]. In this paper, we aim to understand and model this literature gap, opening optimization potential for EA pads to serve as ultra-fast, electrically controllable, planar adhesive clutches.
@@
We developed an electromechanical EA dynamics model, factoring in the dynamics of the drive circuitry, polarization, and mechanical contact between the EA pad and substrate. We test using simulation and experiments how different geometric and operating parameters affect the EA clutch’s engagement and release times to metallic substrates when loaded in shear. We demonstrate EA clutches with engagement times under 15 µs and release times as low as 875 µs, which are 10$\times$ and 17.1$\times$ faster, respectively, than the best times found in prior literature. We anticipate these results can be used to optimize EA clutches for high-bandwidth switchable adhesion applications, such as haptic interfaces, wearable robots, and biomedical devices.
Figure 2: Comparison of engagement and release times of centimeter-scale EA clutches in the literature. Papers that did not report one of the two metrics are placed on the red “Did Not Measure” (DNM) axis. The orange highlighted area indicates sub-millisecond engagement and release times.
II Modeling Electroadhesion
@@
We first modeled the electrical and mechanical dynamics separately to understand the relevant design variables, and then combined them into a full electromechanical dynamics model. While our derivations focus on EA-ICE clutches, the final equations are almost identical for parallel plate EA clutches.
II-A Quasistatic Electrostatic Force
@@
As shown in Fig. 1, we consider an interdigitated comb pattern of $N$ electrodes, each with length $L_{e}$ and spaced $L_{g}$ apart. The EA pad is attracted to a substrate of length $L_{s}$, width $w_{s}$, and thickness $T_{s}$. The dielectric thickness is $T_{d}$, its relative permittivity is $\kappa$, and the average air gap is $T_{air}$. We assume a metal substrate to facilitate modeling, as discussed further in Sec. II-B. Fig. 1(a) shows a diagram of the layer stack’s electric fields and charge densities. After depositing a surface charge density $\sigma$ on an electrode of area $A=L_{e}w_{s}$, a charge $-\sigma$ is induced in the substrate. Inspired by Nakamura and Yamamoto [20], the capacitance between each electrode and the substrate can be modeled as a dielectric and air capacitor in series. The capacitance between $N$ adjacent electrodes (half held at potential $V$, half at ground) is thus:
@@
Figure 3: Comparison of the normal EA force predicted by a finite element simulation against the analytical models in Eq. 4 and Eq. 5.
TABLE I: Electroadhesive clutch and substrate dimensions
@@
$w_{s}\times T_{s}$ (mm×$\times$×mm) 2 $\times$ 2; 2.5 $\times$ 2.5; 3 $\times$ 3
4 $\times$ 2; 5 $\times$ 2.5; 6 $\times$ 3
While the air gap’s importance is well-established empirically in the literature for designing high-force EA clutches [36], Eq. 4 helps explain why. As depicted in Fig. 1(a), the polarized dielectric’s bound charge amplifies the air’s electric field by an extra factor $\kappa$, enabling higher forces for very small air gaps. To verify our model, we developed an electrostatics finite element simulation in Altair Flux for an EA-ICE clutch with nominal dimensions (Table I), a 2 mm wide metal substrate, and a 5 µm air gap. Fig. 3 compares the simulated normal pressure against Eq. 4 and Eq. 5, showing that Eq. 4 well approximates the finite element simulation to within 3%.
While we leave a full retrospection to future work, we propose that factoring in the air gap via Eq. 4 can enable better predictions of EA clutch performance. For example, Hinchet and Shea [12] measured a capacitance of 0.76 $\frac{nF}{cm^{2}}$ for a parallel plate EA clutch at 150 V, which corresponds to an air gap of 610 nm (Eq. 2) and a predicted shear force capacity of 73.9 kPa (Eq. 4). This is much closer to their measured 79.2 kPa than Eq. 5’s predicted 7.2 kPa. Similarly, Berdozzi et al. [37] measured a capacitance change of 76.9 pF for an EA-ICE clutch at 1 kV, corresponding to an air gap of 9.2 µm. Eq. 4 predicts a much closer shear force capacity of 2.54 kPa to their measured 2.02 kPa than Eq. 5’s predicted 15.3 kPa. Diller et al. [6] also noted in their analysis that their data fits $\kappa^{2}$ much better than the $\kappa$ they were using from Eq. 5.
II-B Extending Quasistatic Analysis to the Transient Domain
In this section, we extend the quasistatic analysis in Sec. II-A to the transient domain, considering how polarization dynamics and the high voltage drive circuitry could affect EA dynamics. Contrary to some assertions in prior work [19, 20], we conclude that the upper limit of EA dynamics will be limited largely by mechanical rather than electrical dynamics.
II-B1 Dielectric Constant Dynamics
Sec. II-A’s quasi-static analysis assumed that the substrate’s induced charge instantaneously matches the deposited surface charge $\sigma$. On our time scale, this works for our metal substrate, in which charges can migrate in under a nanosecond. However, dielectrics move charges more slowly as polymer chains and interfacial charges reorient to be in line with the electric field [19].
@@ II-B1 Dielectric Constant Dynamics
where $\delta(t)$ is the Dirac delta function and $E_{\alpha,\beta}(z)$ is the Mittag-Leffler function. Fig. 4 shows the step input response to a fitted Cole-Cole model for P(VDF-TrFE-CFE) based on data from our vendor datasheet and from [17, 18], which tested samples with a similar composition and annealing process to our own. Taking $\kappa_{\infty}=4$ based on [17], the best fit coefficients were $\kappa_{s}=54.2$, $\tau=2.82$ µs, and $\alpha=0.562$. Interpreting the graph, after a charge $\sigma$ is deposited on (or removed from) an electrode, the dielectric can reach 93% polarization (or depolarization) within the first 100 µs. Similar curves can be drawn for other common EA dielectrics, and we conclude that while polarization dynamics set an important upper threshold on EA bandwidth, they are not primarily responsible for the 5-50 ms engagement times seen in prior literature (Fig. 2).
An important note is that the Cole-Cole model applies when the EA pad is driven by a power supply that can both source and sink charge. Prior modeling literature often asserts, based on the language for Maxwell-Wagner interfacial polarization, that the characteristic relaxation time should be on the order of seconds to minutes [19, 20, 21]. However, this assumes the EA pad is disconnected from a power supply and thus can only dissipate charge through internal resistance and stochastic depolarization. Although this is useful in applications such as perching robots, we proceed with the Cole-Cole model since most EA devices always stay connected to their power supply.
@@
II-B2 High Voltage Drive Circuit Dynamics
@@ II-C Mechanical Dynamics
$F_{b}(T_{air})=\left(\frac{96\mu\min(L_{s},w_{s})^{3}\max(L_{s},w_{s})}{\pi^{4}T_{air}^{3}}\right)\dot{T}_{air}$ (10)
where $\mu=$ 18.5 µNs/m2 is air’s dynamic viscosity. To account for the Gaussian distribution of asperity heights, we can find the average damping force by integrating $F_{b}(T_{air})$ with the asperity height’s probability density function [43]:
@@
We can similarly define $F_{ea}^{\prime}$:
@@
where the negative integral range accounts for when the dielectric compresses against asperities. These alterations have very little effect when $T_{air}>>\sigma_{d}$, but for our nominal $T_{d}$ it can increase $F_{ea}$ by up to 11.1$\times$ when $T_{air}\approx\sigma_{d}$.
II-D Full Electromechanical Dynamics Model
Combining the equations from the last three sections together, we can write equations of motion in the $z$ axis for the dielectric and substrate. We assume the substrate is much heavier than the dielectric (i.e., fixed in z-height), while the dielectric is free to move. The dielectric also has a small preload force $F_{preload}$ applied to help prevent wrinkles and ensure even contact between the surfaces. The dielectric’s equation of motion along the $z$ axis is:
$m_{d}\ddot{T}_{air}=F_{k}-F_{b}^{\prime}-F_{ea}^{\prime}-m_{d}g-F_{preload}$ (13)
where the first term on the right side comes from contact mechanics (Eq. 8), the second term comes from damping (Eq. 11), the third comes from electrostatics (Eq. 12, but incorporating the rise and fall times of $\kappa$ and $V$ as described in Sec. II-B), and the fourth comes from gravity on the dielectric mass $m_{d}$. The system of equations can be solved for as an initial value problem with the state $[T_{air},\dot{T}_{air}]$.

### v2

@@ Abstract
Electroadhesive clutches are electrically controllable switchable adhesives commonly used in soft robots and haptic user interfaces. They can form strong bonds to a wide variety of surfaces at low power consumption. However, electroadhesive clutches in the literature engage to and release from substrates several orders of magnitude slower than a traditional electrostatic model would predict. Large release times, in particular, can limit electroadhesion’s usefulness in high-bandwidth applications. We develop a novel electromechanical model for electroadhesion, factoring in polarization dynamics, the drive circuitry’s rise and fall times, and contact mechanics between the dielectric and substrate. We show in simulation and experimentally how different design parameters affect the engagement and release times of centimeter-scale electroadhesive clutches to metallic substrates, and we find that the model accurately captures the magnitude and trends of our experimental results. In particular, we find that higher drive frequencies, narrower substrate aspect ratios, and faster drive circuitry output stages enable significantly faster release times. The fastest clutches have engagement times less than 15 µs and release times less than 875 µs, which are 10$\times$ and 17.1$\times$ faster, respectively, than the best times found in prior literature on centimeter-scale electroadhesive clutches.
keywords:
Electroadhesion , switchable adhesion , dynamics , modeling , soft actuators
journal: Extreme Mechanics Letters
[…]
organization=Department of Mechanical Engineering at Stanford University, city=Stanford, state=CA 94305, country=USA
1 Introduction
@@
Switchable adhesives can dynamically adhere to and release from contacting surfaces. They are widely used in soft robots, haptic interfaces, and biomedical devices, with applications in robotic grippers [1, 2], wall climbing and perching robots [3, 4, 5], exoskeletons [6, 7], haptic user interfaces [8, 9, 10, 11], and kinesthetic garments [12, 13]. Different switchable adhesion mechanisms can be compared by their shear force capacity when gripping different substrates and by their engagement and release speeds [14]. Mechanisms with high shear force capacities, like granular jamming actuators and shape memory polymers, often require long switching times due to viscoelastic effects, while faster adhesives like liquid crystal elastomers and gecko-inspired adhesives often have caveats like low force capacity or poor adhesion to rough surfaces [15, 16].
@@
Electroadhesion (EA), or the electrostatic attraction between two contacting materials separated by a dielectric and held at a potential difference, has emerged as an effective electrically triggered switchable adhesive. Its scalable manufacturing process, high shear force capacity against a wide range of substrates, relatively fast actuation speeds, low power consumption, and low profile have made EA useful for space- and power-constrained applications. In this paper, we focus on a common implementation, shown in Fig. 1(a), where one of the contacting materials is a dielectric patterned with interdigitated comb electrodes while the other contacting material can be any conductive or insulating substrate. This form factor, originally developed in 1968 by Krape [17], sacrifices some adhesive force capacity compared to a parallel plate configuration in which both contacting materials have electrodes. However, EA with Interdigitated Co-planar Electrodes (shortened to EA-ICE in this paper) allow all the electronics to be concentrated into a single surface, reducing bulk and enabling the clutch to adhere to small, delicate objects and generic surfaces in the field.
@@
Figure 1: (a) Schematic diagram and (b) picture of an electroadhesive clutch. When a voltage $V$ is applied, the polarization of bound charges in the dielectric amplifies the field strength, attracting the film to the metal substrate. The resulting friction can counteract applied shear loads on the substrate. (c) Comparison of engagement and release times of centimeter-scale EA clutches in the literature. Papers that did not report one of the two metrics are placed on the red “Did Not Measure” (DNM) axis. The orange highlighted area indicates sub-millisecond engagement and release times. The raw data is provided in Supplementary Materials, Table S1. [1, 2, 6, 7, 9, 13, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]
@@
Despite electroadhesion’s widespread utility, there is a lack of accurate dynamics models for how quickly EA clutches can adhere to and release from surfaces. Prior work has shown that the dielectrics themselves can be charged and discharged on the order of microseconds [32, 33, 34], and sub-millimeter-scale electroadhesive systems can indeed measure engagement and release times on this order of magnitude [35, 36, 37, 38]. However, as shown in Fig. 1(c), every paper we found on centimeter-scale electroadhesive clutches reported engagement times from hundreds of microseconds to milliseconds and release times ranging from tens of milliseconds to minutes. While some models predicting minute-scale release times have been proposed (although, to our knowledge, never compared numerically against experimental data) [18, 19, 39, 40], experimental results in the microsecond to millisecond range have largely gone unexplained. Slow release times limit EA’s scalability for high-bandwidth applications [11, 41]. Slow dynamics also leads to “stick-slip” behavior, where after an applied load exceeds an EA clutch’s force capacity the clutch slips, losing up to 70% of its frictional capacity and requiring tens of milliseconds to re-engage [7, 20, 42]. In this paper, we aim to understand and model this literature gap, opening optimization potential for EA pads to serve as ultra-fast, electrically controllable, planar adhesive clutches.
@@
We developed an electromechanical EA dynamics model, factoring in the dynamics of the drive circuitry, polarization, and mechanical contact between the EA pad and substrate. These considerations enable our model, unlike prior EA dynamics models, to predict microsecond- to millisecond-scale engagement times and millisecond-scale release times in line with the range seen in previous experimental studies. We test using simulation and experiments how different geometric and operating parameters affect the EA clutch’s engagement and release times to metallic substrates when loaded in shear. The fastest clutches are demonstrated to have engagement times less than 15 µs and release times less than 875 µs, which are 10$\times$ and 17.1$\times$ faster, respectively, than the best times found in prior literature on centimeter-scale EA clutches. These improvements, particularly to release time, enable our EA clutches’ frequency responses to have -3 dB bandwidths over 10 kHz, and we also show that EA clutches tested in conditions that result in better release times also often recover faster after a slip. We anticipate these results can be used to optimize EA clutches for high-bandwidth switchable adhesion applications, such as haptic interfaces, wearable robots, and biomedical devices.
2 Modeling Electroadhesion
@@
We first modeled the electrical and mechanical dynamics separately to understand the relevant design variables, and then combined them into a full electromechanical dynamics model. While our derivations focus on EA-ICE clutches adhering to metallic substrates, the final equations are almost identical for parallel plate EA clutches, as shown in Supplementary Materials, Sec. S2.
2.1 Quasistatic Electroadhesive Force
@@
As shown in Fig. 1(a, b), we consider an interdigitated comb pattern of $N$ electrodes, each with length $L_{e}$ and spaced $L_{g}$ apart. The EA pad is attracted to a substrate with overlap length $L_{s}$, width $w_{s}$, and thickness $T_{s}$. The dielectric thickness is $T_{d}$, its relative permittivity is $\kappa$, and the average air gap is $T_{air}$. We assume a metal substrate to facilitate modeling, as discussed further in Sec. 2.2. Fig. 1(a) shows a diagram of the layer stack’s electric fields and charge densities. After depositing a surface charge density $\sigma$ on an electrode with overlap area $A=L_{e}w_{s}$, a charge $-\sigma$ is induced in the substrate. Inspired by Nakamura and Yamamoto [19], the capacitance between each electrode and the substrate can be modeled as a dielectric and air capacitor in series. Assuming for now that both the electrode and substrate surfaces are perfectly flat, the capacitance between $N$ adjacent electrodes (half held at potential $V$, half at ground) is thus:
@@
[…]
Figure 2: (a) Schematic of the geometry used for finite element simulation. (b) Comparison of the normal EA force predicted by a finite element simulation against the analytical models in Equations 4 and 5 for dielectrics with relative permittivities $\kappa=3$ and $\kappa=50$. (c, d) Finite element analysis of the electric field (colored arrows) and potential (grayscale iso-contours) through an electroadhesive layer stack for a metallic substrate. The dielectric is a 24 µm thick film dielectric with (c) $\kappa=3$ and (d) $\kappa=50$, and the air gap depicted here is 5 µm thick. Both (c) and (d) use the same potential and electric field color scheme.
Table 1: Electroadhesive clutch and substrate dimensions, as shown in Fig. 1(a, b)
@@
$w_{s}\times T_{s}$ (mm×$\times$×mm) 2 $\times$ 2; 2.5 $\times$ 2.5; 3 $\times$ 3; 4 $\times$ 2; 5 $\times$ 2.5; 6 $\times$ 3
While we leave a full retrospection to future work, we propose that factoring in the air gap via Eq. 4 can enable better predictions of EA clutch performance. For example, Hinchet and Shea [13] measured a capacitance of 0.76 nF cm-2 for a parallel plate EA clutch at 150 V, which corresponds to an air gap of 610 nm (Eq. 2) and a predicted shear force capacity of 73.9 kPa (Eq. 4). This is much closer to their measured 79.2 kPa than Eq. 5’s predicted 7.2 kPa. Similarly, Berdozzi et al. [45] measured a capacitance change of 76.9 pF for an EA-ICE clutch at 1 kV, corresponding to an air gap of 9.2 µm. Eq. 4 predicts a much closer shear force capacity of 2.54 kPa to their measured 2.02 kPa than Eq. 5’s predicted 15.3 kPa. Diller et al. [7] also noted in their analysis that their data fits $\kappa^{2}$ much better than the $\kappa$ they were using from Eq. 5.
As a technical note, our derivation above only considers Coulombic electroadhesion, which affects both insulating and semiconducting dielectrics and which can create electrostatic attractive forces between both contacting and non-contacting surfaces. We do not extend our model to Johnsen-Rahbek electroadhesion, which primarily only affects semiconducting dielectrics in close contact with the substrate, because the dielectrics in this study are insulators ($\rho>10^{10}$ $\Omega$ cm). We refer the reader to [19, 46] for an in-depth comparison between Coulombic and Johnsen-Rahbek electroadhesion.
2.2 Extending Quasistatic Analysis to the Transient Domain
In this section, we extend the quasistatic analysis in Sec. 2.1 to the transient domain, considering how polarization dynamics and the high voltage drive circuitry could affect EA dynamics. Contrary to assumptions in prior modeling work [18, 19], we conclude that the upper limit of EA dynamics will be limited largely by mechanical rather than electrical dynamics.
2.2.1 Dielectric Constant Dynamics
Sec. 2.1’s quasi-static analysis assumed that the substrate’s induced charge instantaneously matches the deposited surface charge $\sigma$. This assumption is valid at our time scale for a metal substrate, in which charges can migrate in under a nanosecond. However, dielectrics move charges more slowly as polymer chains and interfacial charges reorient to be in line with the electric field [18].
@@ 2.2.1 Dielectric Constant Dynamics
where $\delta(t)$ is the Dirac delta function and $E_{\alpha,\beta}(z)$ is the Mittag-Leffler function. Fig. 3 shows the step input response to a fitted Cole-Cole model for P(VDF-TrFE-CFE) based on data from our vendor datasheet and from [32, 33], which tested samples with a similar composition and annealing process to our own. Taking $\kappa_{\infty}=4$ based on [32], the best fit coefficients were $\kappa_{s}=54.2$, $\tau=2.82$ µs, and $\alpha=0.562$. Interpreting the graph, after a charge $\sigma$ is deposited on (or removed from) an electrode, the dielectric can reach 93% polarization (or depolarization) within the first 100 µs. Similar curves can be drawn for other common EA dielectrics, and we conclude that while polarization dynamics set an important upper threshold on EA bandwidth, they are not primarily responsible for the 5-50 ms engagement times seen in prior literature (Fig. 1(c)).
An important note is that the Cole-Cole model applies when the EA pad is driven by a power supply that can both source and sink charge. Prior modeling literature often asserts, based on the language for Maxwell-Wagner interfacial polarization, that the characteristic relaxation time should be on the order of seconds to minutes [18, 19, 39]. However, this assumes the EA pad is disconnected from a power supply and thus can only dissipate charge through internal resistance and stochastic depolarization. Although this is useful in applications such as perching robots or low-bandwidth tactile displays, where EA pads adhere to an insulating substrate using a power supply that intentionally does not sink the interfacial polarization charge [3, 34], we proceed with the Cole-Cole model since most EA devices always stay connected to their power supply. Additional discussion and finite element analyses supporting a Cole-Cole polarization model instead of a Maxwell-Wagner polarization model for our application are presented in Supplementary Materials, Sec. S3.
@@
2.2.2 High Voltage Drive Circuit Dynamics
@@ 2.3 Mechanical Dynamics
$F_{b}(T_{air})=-\left(\frac{96\mu_{air}\min(L_{s},w_{s})^{3}\max(L_{s},w_{s})}{\pi^{4}T_{air}^{3}}\right)\dot{T}_{air}$ (10)
where $\mu_{air}=$ 1.85 $\cdot 10^{-5}$ N s m-2 is air’s dynamic viscosity. To account for the Gaussian distribution of asperity heights, we can find the average damping force by integrating $F_{b}(T_{air})$ with the asperity height’s probability density function [46, 52]:
@@
Conveniently, the average gap between surfaces used in contact mechanics models is the same $T_{air}$ used in our EA normal force equation $F_{ea}(T_{air})$ (Eq. 4). Thus, to account for the effect of surface roughness, we can use a similar integral over the asperity height probability density function to define the average EA normal force $F_{ea}^{\prime}$:
@@
where the negative integral range accounts for when the dielectric compresses against asperities. These alterations have very little effect when $T_{air}\gg\sigma_{d}$, but for our nominal $T_{d}$ it can increase $F_{ea}^{\prime}$ by up to 11.1$\times$ when $T_{air}\approx\sigma_{d}$. This formulation only accounts for EA force in the $z$ direction and assumes the influence of fringing fields between asperities is negligible, similar to prior work [46, 52], although future work should further investigate the validity of this assumption.
We also explored a similar averaging formulation of the parallel plate capacitance equation in Eq. 1, but we found it made very little difference for our system. This matches the findings of prior work on capacitive contact profilometers [53, 54, 55, 56], which generally find that micrometer-scale surface roughnesses only alter the ideal flat plate capacitance by a few percent over centimeter-scale overlap areas so long as $T_{air}>\sigma_{d}$. Using similar reasoning, prior electroadhesion studies have primarily used parallel plate capacitance models to estimate their systems’ average air gap [11, 34, 57, 58].
2.4 Full Electromechanical Dynamics Model
Figure 4: Free body diagram of forces on the dielectric and substrate.
Combining the equations from the last three sections together, we can write equations of motion in the $z$ axis for the dielectric and substrate. We assume the substrate is much heavier than the dielectric (i.e., fixed in $z$-height), while the dielectric is free to move in the $z$ direction. The dielectric also has a small preload force $F_{preload}$ applied to help prevent wrinkles and ensure even contact between the surfaces, in line with prior experimental work [42, 59, 60, 61]. Referencing the free body diagram in Fig. 4, the dielectric’s equation of motion along the $z$ axis is:
$m_{d}\ddot{T}_{air}=F_{k}+F_{b}^{\prime}-F_{ea}^{\prime}-m_{d}g-F_{preload}$ (13)
where the first term on the right side comes from contact mechanics (Eq. 8), the second term comes from damping (Eq. 11), the third comes from electrostatics (Eq. 12, but incorporating the rise and fall times of $\kappa$ and $V$ as described in Sec. 2.2), and the fourth comes from gravity on the dielectric mass $m_{d}$.